Bridgeport Company purchased a delivery truck for $27,000 on January 1, 2020. The truck has an expected salvage value of $1,500, and is expected to be driven 102,000 miles over its estimated useful life of 10 years. Actual miles driven were 12,800 in 2020 and 14,300 in 2021.

Compute depreciation expense for 2020 and 2021 using (1) the straight-line method, (2) the units-of-activity method, and (3) the double-declining-balance method. (Round depreciation cost per unit to 2 decimal places, e.g. 0.50 and depreciation rate to 0 decimal places, e.g. 15%. Round final answers to 0 decimal places, e.g. 2,125.)

Depreciation Expense

2020

2021

(1) Straight-line method $

$

(2) Units-of-activity method $

$

(3) Declining-balance method $

$
Answer #1
1
Straight-line method
Depreciation Expense:
2020 2550 =(27000-1500)/10
2021 2550 =(27000-1500)/10
2
Units-of-activity method:
Per mile depreciation 0.25 =(27000-1500)/102000
Depreciation Expense:
2020 3200 =12800*0.25
2021 3575 =14300*0.25
3
Declining-balance method:
Depreciation rate 20% =1/10*2
Depreciation Expense:
2020 5400 =27000*20%
2021 4320 =(27000-5400)*20%